OpenStudy (anonymous):

evaluate each infinite geometric series 3+2+(4/3)+(8/9)

OpenStudy (apoorvk):

you're basically multiplying each term by 2/3 - so the common ratio 'r' is 2/3 now for a GP, where the common ration is between 0 and 1, the sum upto infinite term is a finite no. S = a/(1-r) S = sum a=first term r = common ratio. Evaluate. can you?

OpenStudy (anonymous):

2/3 (4/3)/2=2/3 (8/9)/(4/3) = 2/3 therefore, it is a G.S. with common ratio = 2/3 Sum of G.S.= a(1-r^n) / (1-r) =(3)(1-(2/3)^4) /(1-2/3) =9(1-(2/3)^4) =.....

OpenStudy (anonymous):

that will give me the answer?

OpenStudy (anonymous):

65/9

OpenStudy (apoorvk):

How we arrive at "S = a(1-r)" is this way Sum of any GP upto 'n' terms is: \[S = \frac {a(1-r^n)}{(1-r)}\] When you put in n=infinity in "r^n", and r is a positive fraction less than 1, r multiplied by itself infinite times will render it so small, that r^n can almost be neglected. So, you are left with: S = a/(1-r)